# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

### v6

@@ 1 Introduction
To achieve this, as shown in Figure 1, Data-Copilot contains three steps for data exploration:
@@
    Request Exploration: Data-Copilot explores potential requests and mines a broader range of human need scenarios from data. It involves a "self-instruct" process to generate massive simulated requests based on all data schemas (types and formats) and seed requests.
@@
    Interface Design: Our agent generates code for each request and then considers whether these requests have similar demands and whether general interfaces can be abstracted.
@@
    Interface Optimization: To improve the quality of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces by reflecting on feedback from the compiler.
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different request scenarios, Data-Copilot can flexibly design different interface invocation procedures, such as serial workflows, i.e., step-by-step invocation, parallel workflows, and loop workflows. Our contributions are threefold:
Figure 2: Overview of Data-Copilot. Data Exploration: It devises a self-request process to generate numerous potential requests based on data schemas. Then it interleaves to design and optimize the interfaces to solve these requests. The whole exploration-design-optimization process is operated in advance, generating many general and error-free interface modules, such as data retrieval, computation, and visualization. Workflow Deployment: Upon receiving real-time requests, LLM invokes these interfaces, forming an interface workflow with different structures for request solving.
@@ 2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-request process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these "self-instruct" requests, it generates code and abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and reflects on the code compiler feedback, ensuring each interface is universal and correct (Step 3: optimization). Finally, Data-Copilot acquires many generic, error-free interfaces for subsequent use. The whole exploration-design-optimization process is operated in advance.
When faced with real-time requests (Section 2.2), Data-Copilot adeptly comprehends user intentions, and invokes pre-designed interfaces for request solving, significantly reducing the need for tedious labor. We provide a detailed prompt for two phases in Section F.1 and F.2.
@@ 2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to humans, Data-Copilot aims to explore data and derive insights from vast data sources. These insights include understanding the relations between the data, discovering the potential requests associated with the data, and organizing the code framework efficiently. Based on this, Data-Copilot summarizes these exploratory insights into many reusable code components and optimizes its performance. It significantly reduces the code complexity for real-time responses, thus improving the success rate.
Request Exploration
To explore data and mine insights, we design a self-request process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs must generate requests specifically related to these data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.
@@
###Parsing file for Stock_data:{...},
How to ensure the quality of the generated requests? A common issue is that the LLM often proposes a request about the data that doesn’t exist. To address this, we design a backward verification process to check the rationality of these requests. Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests. Upon manually check some requests, we find that the generated requests generally met our expectations.
@@
After generating massive requests through the self-request, Data-Copilot considers abstracting universal interfaces from these requests. First, we need to define what an interface is in our paper. Similar to human programming practices, we refer to a function module abstracted from the code as an interface. Each interface is composed of a name, parameters, a functional description, and the implementation code, performing specific functions such as data retrieval, computing, and plotting. To abstract general interfaces from all requests, we design an interwoven process, performing interface design and interface optimization alternately for each request.
Firstly, we feed related data parsing files and existing interfaces stored in the interface library (empty at first iteration) into LLMs as a prompt. LLMs need to generate a complete code for each request. When generating, LLMs are instructed to prioritize using the existing interfaces from the previous iterations. If these interfaces are insufficient to address the request, the LLMs design a new interface. As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...). As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
@@
To improve the quality of interfaces, Data-Copilot optimizes them in two ways: merging similar interfaces to make them more universal and reflecting on compiler feedback to reduce errors. This optimization process is intertwined with the interface design for each request.
Firstly, similar to software developers wrapping similar code modules to reduce redundancy, Data-Copilot considers whether each newly designed interface can be merged with the previous interfaces. LLMs are instructed to check the functionality, parameters, and logic of the new and existing interfaces, and merge those similar interface modules to create a new one. As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...). This interface optimization process makes each interface more general and rational.
Besides, to reduce code hallucinations or formatting issues, Data-Copilot also self-reflects on the feedback of code execution for each request if compilation errors occur. We have LLMs perform self-reflection until the code passes or the maximum retry number is reached. This process ensures the correctness and usability of each interface.
For each request, Data-Copilot alternates the above two steps: interface design and optimization until all the requests can be satisfied by these interfaces. Finally, as shown in Figure G3, Data-Copilot produces five types of interfaces: data acquisition, processing, prediction, visualization, and DataFrame manipulation. We provide some detailed cases in Section G.1. Besides, to handle newly added data sources, Data-Copilot can repeat the aforementioned exploration-design-optimization process on new data. It can create some new interface modules for these new data.
@@ 2.2 Workflow Deployment
In the first phase, Data-Copilot designs many generic interfaces with different functionalities. They are versatile and easy to use, simplifying the code generation of real-time responses. As shown in Figure 2, when faced with human requests, Data-Copilot only needs to invoke related interfaces and form an interface workflow. A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex call flow diagram. This workflow transforms raw data into visualized results. The detailed prompts and cases are shown in Sections F.2 and G.2.

### v7

@@ 1 Introduction
As shown in Figure 1, Data-Copilot contains three key steps for data exploration:
@@
    Explore data and synthesize requests: Data-Copilot discovers potential requests and a broader range of human need scenarios from data. It involves a "self-exploration" process to generate massive synthesized requests based on all data schemas and seed requests.
@@
    Interface design and test: Data-Copilot generates modular code for request solving. Each interface is a functional module, and test cases are generated to verify the designed interface.
@@
    Interface optimization: To improve the versatility of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces based on the compiler’s feedback.
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop. It can even output a hybrid form of interface workflows and raw code for some "unfamiliar" requests. Our contributions are threefold:
Figure 2: Overview of Data-Copilot. Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources. Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces one by one. This exploration-design-optimization process is operated in advance. Workflow Deployment: Upon receiving real-time requests, Data-Copilot can quickly invoke existing interfaces and deploy a workflow for familiar requests, or flexibly combine interfaces and raw code for "uncovered" requests.
@@ 2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these synthesized requests, it abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization). The whole exploration-design-optimization process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request. In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed. We provide a detailed prompt for two phases in Section F.1 and F.2.
@@ 2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data. Then Data-Copilot abstracts these exploratory insights into numerous reusable code components (interfaces), testing their correctness and optimizing their generality. This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
Self-Exploration Request
To explore data and mine insights, we design a self-exploration process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs should generate requests specifically based on provided data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.
@@
###Parsing file for Stock_data:{...}, ....
How to ensure the quality of the generated requests? A common issue is that the LLM often proposes a request about the data that doesn’t exist. To address this, we design a backward verification process to check the rationality of these requests. Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests. Upon manually checking some requests, we found that the generated requests generally met our expectations.
@@
After generating massive requests through self-exploration, Data-Copilot abstracts many universal interfaces from these requests. First, we have to clarify what an interface is in our paper. Similar to human-defined functions, an interface is a code module consisting of a name, parameters, a functional description, and an implementation code. It performs specific tasks such as data retrieval, computation, and visualization. In our paper, each interface is designed, tested, and optimized iteratively by Data-Copilot.
First, starting from the initial request, we iteratively feed self-exploration requests and related data parsing files into LLMs, prompting the LLM to design complete code modules for request solving. Each code module corresponds to an interface and is stored in the interface library. During each iteration, LLMs are instructed to prioritize utilizing the existing interfaces within the library. If the available interfaces are insufficient for request solving, the LLMs design a new interface.
As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...). As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
Interface Testing
After designing a new interface for a request (request-i $\rightarrow$ interface-i), Data-Copilot autonomously tests its correctness based on compiler feedback. First, Data-Copilot uses the request-i as the seed to generate massive similar requests as test cases. Then it tests the new interface-i one by one. E.g., in Fig. 2, for newly designed interface-i: Query-GDP(), Data-Copilot mimics request-1 to generate many similar requests to test the Query-GDP(): "I want to see USA’s GDP over the past 10 years", "I want to see China’s GDP for the last year". If the interface passes all test cases, it is retained; otherwise, Data-Copilot utilizes compiler error feedback for self-reflection, correcting the erroneous code snippets until it successfully passes the tests.
@@
To further improve the generality of interfaces, Data-Copilot also optimizes similar interfaces. Similar to how human software developers, each time a new interface is designed, this optimization process is triggered: evaluating whether the newly designed interface can be merged with previous ones.
  • Retrieve similar interfaces. After a new interface is designed, we first retrieve the top N similar interfaces from the existing interface library. Specifically, we use gte-Qwen1.5-7B-instruct (Li et al., 2023d) to obtain embeddings of each interface code and then calculate their similarity.
  • Decide whether to merge. Then LLMs are prompted to evaluate the functionality, parameters, and logic of the new interface and the top N similar interfaces, autonomously deciding whether to merge the new interface with existing ones. If the LLM deems no merging necessary, the new interface is retained. As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...). This optimization process makes each interface more general and unique, reducing interface redundancy.
  • Test the optimized interface. When two interfaces are merged, Data-Copilot also needs to test the newly merged interface. The test cases from the two original interfaces are used to validate the merged interface. The output of the new interface must be consistent with the two original interfaces in both format and content.
For each request, Data-Copilot alternates the above three steps: interface design, testing, and optimization until all the requests can be covered by these interfaces. As shown in Figure G3, Data-Copilot designs many interfaces for different task and operation types, e.g., data acquisition, prediction, visualization, and DataFrame manipulation. We provide some detailed cases in Section G.1.
@@ 2.2 Workflow Deployment
In the first phase, Data-Copilot designs many generic interfaces with different functionalities in advance. They are versatile and error-free, capable of handling the majority of human requests in real-world scenarios. As shown in Figure 2, Data-Copilot only needs to invoke the relevant interfaces for human requests, which is more accurate and efficient (interface workflow). Furthermore, for some long-tail or new requests that cannot be covered by the existing interfaces, Data-Copilot can flexibly generate raw code to handle them (interface-code hybrid strategy) and document these requests for updating the interface library in the future.
Task Selection and Interfaces Retrieval
Considering the significant differences between the interfaces involved in different requests, it is unnecessary to load all interfaces every time. We design a simple yet efficient interface retrieval strategy: hierarchical retrieval. Specifically, we organize all designed interfaces in a hierarchical structure. Each interface is grouped into different tasks (stock task, fund task, etc.) and different operation types (data acquisition, processing, visualization, etc.). Upon receiving a real-time request, Data-Copilot first determines the appropriate task types and required operation types and then loads the interfaces associated with these types for subsequent workflow planning. This can significantly reduce the number of interfaces in the prompt.
Planning Workflow with Different Structures
After reading the descriptions of all retrieved interfaces, Data-Copilot plans an interface workflow to process the user request. A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex loop structures. Therefore, Data-Copilot needs to autonomously determine which interfaces to invoke, in what sequence, and which parameters to populate. The interface invocation workflow is output in JSON format: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}. The prompts and cases are shown in Sections F.2 and G.2.
As shown in Figure 2, Data-Copilot designs interface workflows with different structures, e.g., sequential, parallel, or loop. For instance, the first request in Figure 2: "Compare the return of CSI-300, GEM and CSI-1000 this year", Data-Copilot plans a parallel workflow with 5 steps: Data Acquisition(), Calcuate Return() in parallel for calculating three indices, followed by Visualization(). In the second case, it deploys a loop workflow by calling the Loop().
Interface-Code Hybrid Generation
In the real world, it is inevitable to encounter "uncovered" requests that cannot be addressed by existing interfaces. As a remedy, Data-Copilot adopts an interface-code hybrid generation strategy. Specifically, Data-Copilot prioritizes invoking existing interfaces to resolve user requests. It can solve most common user requests. However, if the deployed workflow continues to fail, or if Data-Copilot proactively determines that the current request cannot be resolved by existing interfaces, our system is allowed to generate raw code directly or generate a combination of raw code and interfaces for request solving. This design endows Data-Copilot with the flexibility to handle diverse requests.
Besides, each time Data-Copilot encounters such "uncovered" requests, it also records them in the document. After accumulating a certain number of new requests, we re-initiated the Data Exploration stage (Section 2.1) to develop new interfaces for these newly emerged requests. It empowers Data-Copilot to continuously update its interface library throughout human interactions.
Table 1: Statistics on human-proposed requests and self-exploration dataset.
Source #Cases Form Usage Ratio of 4 Tasks Ratio of 3 Complexities
Human-proposed 173 Query Seed Set 8: 4: 4: 1 2.0: 1.5: 1
Self-Exploration 3000 Query Interface Design 8.7: 3.3: 4.2: 1 1.9: 1.5: 1
Self-Exploration & Human 547 Query, Label Test Set 8.6: 3.3: 4.3: 1 1.8: 1.4: 1
3 Qualitative Evaluation
3.1 Environment and Data Sources
Data-Copilot is evaluating on Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. Similar to many real-world financial companies, Data-Copilot utilizes data interfaces provided by Tushare [^1] https://tushare.pro/ to access vast amounts of financial data, including time-series data spanning over 20 years for more than 4,000 stocks, funds, and futures. In the first phase, we use a strong LLM (e.g., gpt-4-turbo) for data exploration and a lightweight LLM (e.g., gpt-3.5-turbo) for workflow deployment.
3.2 The Creation of Dataset
Human-proposed Request as Seed. We invite 10 students in economics to submit 50 requests each. These requests cover a wide range of common needs, including stock tasks, fund tasks, and corporate tasks, and also involve different difficulty levels. We treat these 500 human-proposed requests as a distribution of real-world scenarios, where the ratio of four task types is Stock: Fund: Corporation: Other = 8: 4: 4: 1. Then we filter out highly similar requests by computing their sentence embedding. Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
Self-Exploration Request Set. As mentioned in Section 2.1, we employ self-exploration to expand the request set. Specifically, we feed 173 human-proposed seed requests along with all data descriptions and schema into GPT-4 for data exploration and request synthesis. When self-exploration, we use keywords to control the distribution of the synthesized requests, ensuring that they can closely align with real-world scenarios. In this phase, we generate 6480 requests. After that, we filter out highly similar requests and retain 3547 requests. Then we adopted a stratified sampling strategy to sample 547 instances as the test set from each task type. The remaining 3000 requests are used for interface design. We analyze the number of subsets categorized by four task types (stock, fund, corporation, and others) and three complexity levels (single-entity, multi-entities, and multi-entities with complex relations). The results indicate that our synthesized requests closely align with human distribution. The detailed statistics of three datasets are shown in Tables 1, B1 and B1.
Table 2: Human evaluation on the human-proposed seed set and self-exploration requests.
Number Task Difficulty Request Rationality Expression Ambiguity Answer Accuracy
Seed Set 173 3.5 4.3 4.4 -
Test Set 547 3.9 4.2 4.0 4.8
Annotate Answer Table for Test Set. Human annotators are instructed to annotate data-table as answers for 547 test requests, which is used for benchmarking our system’s output. Specifically, annotators manually retrieve and process the corresponding data based on the testing request, and record the final data table before chart plotting. These data tables are used as answer annotations. For example, for the request: "I want to see China’s GDP trend over the past 5 years", the answer data-table is "[2023: 17.8 trillion, 2022: 17.9 trillion, 2021: 17.8 trillion, 2020: 14.6 trillion, 2019: 14.3 trillion]".
The Quality of Self-exploration Requests. To assess the quality of self-exploration requests, we invited four additional graduate students to manually evaluate our test set (546 requests) and human-proposed requests (173 seed set) on four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy. We provide detailed guidance for each criterion in Section B.2. The results of two sets are shown in Table 2. We observed that the synthesized requests exhibit a comparable quality to human-proposed requests, with slightly higher difficulty, ambiguity, and similar rationality. It ensures that our test set can reflect most of the real-world demands. Besides, the evaluators gave a high score of 4.8 on the label (answer table) of our test set, which also ensures the accuracy of our dataset.
3.3 Qualitative Results
Data-Copilot designs versatile interfaces through data exploration. After exploring the Chinese financial data and processing 3000 self-exploration requests, Data-Copilot designed a total of 73 interfaces, comprising 5 functionalities and 4 tasks: data acquisition, index calculation, table manipulation, visualization, and general processing logic. These tasks include stocks, funds, corporations, and others. We present some representative interfaces and the statistics of the interfaces in Figure G3. Most interfaces are highly versatile with multiple functions. For example, get-stock-prices-data() interface can retrieve data on a daily, weekly, or monthly basis. These versatile interfaces endow the Data-Copilot with the capability to address complex requests.
Data-Copilot can flexibly deploy interfaces or raw code for diverse demands. As mentioned in  Section 2.2, for some "new" requests, Data-Copilot is capable of addressing them by combining both raw code and existing interfaces. As shown in Figure 2 (3rd example), the user requests a radar chart, which was not directly supported by the available interface. Data-Copilot first leverages the existing interface to retrieve and process the data. Then based on the dataframe produced by the last interface, it generates raw code to render a radar chart. This design grants Data-Copilot greater flexibility to handle the constantly emerging new requests.
Interface Workflow is human-friendly, offering higher interpretability. Upon executing the workflow via the compiler, Data-Copilot produces the visualized results. Also, it provides a comprehensive summary of the entire processing process based on the workflow. As shown in Figures 2, G2, G4, G6, G5 and G7, Data-Copilot’s outputs (workflow, visualized results and summary) are highly intuitive and clear. It encapsulates complex code processing logic into a step-by-step interface invocation workflow, which is easier for humans to understand and inspect.
Table 3: Pass@1 of Data-Copilot and direct code generation (Raw Code) when using different LLMs for workflow deployment: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Llama-3-70B-chat, Llama-2-70B-chat, Llama-2-13B-chat, Codellama-13B, Vicuna-13B-v1.5).
GPT-4o GPT-4 GPT-3.5 Llama3-70B Llama2-70B Llama2-13B Codellama13B Vicuna
Raw Code 73.6 70.0 28.5 52.2 29.6 17.1 21.0 17.3
Ours 77.9 $\uparrow$4.3 74.6 $\uparrow$4.6 70.2 $\uparrow$41.7 70.9 $\uparrow$18.7 49.3 $\uparrow$19.7 36.9 $\uparrow$19.8 50.8 $\uparrow$29.2 33.2$\uparrow$15
4 Quantitative Evaluation
4.1 Experiments Settings
Baselines and Experiments Details We first compare Data-Copilot with direct code generation (denoted as Direct-Code) using different LLMs. Besides, we also compare Data-Copilot with many prompting and agent strategies, including: 1. ReAct (Yao et al., 2022): LLMs use neural language (thought) and code to resolve the user request step-by-step. 2. Reflexion (Shinn et al., 2023): LLM can reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 3. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. For a fair comparison, Data-Copilot employs the same LLM as all baselines for evaluation when invoking pre-designed interfaces. The detailed prompts of all baselines are shown in Section F.3.
Evaluation Methods. All methods are evaluated on the 547 testing instances. We focus on the accuracy of the final results and the efficiency of the response. Considering that the output of most requests is images or charts, we evaluate from two perspectives: whether the data used for plotting is correct and whether the drawn image meets human requirements. We use GPT-4-turbo and GPT-4o to conduct automatic evaluations. (1) Data-table Evaluation: We first obtain the dataframe before visualization, then use GPT-4-turbo to compare this data table with the answer table (labeled by humans in Section 3.2). (2) Image Evaluation: To evaluate whether the final image meets human requests, we feed the human request, human-annotated answer table, and the output image into GPT-4o for evaluation. We instruct GPT-4o to check whether visual elements in the image meet the human requests, including numerical points, lines, axes, image aesthetics, and style. Finally, the results of data and image evaluation are combined as Pass@1. Lastly, we calculate the total number of tokens consumed by each method (denoted as #Token), which represents the response efficiency.
Table 4: We demonstrate the accuracy and efficiency of all methods with gpt-3.5-turbo
Methods Pass@1(%) #Token
Direct-Code 28.5 $\pm$2.1 823
ReAct 44.1 $\pm$3.4 1515
Reflexion 59.0 $\pm$3.9 2463
Multi-Agent 57.4 $\pm$1.8 2835
Data-Copilot 70.2 $\pm$2.5 561.2
Data-Copilot + ReAct 71.5 $\pm$1.7 834
Data-Copilot + Reflexion 71.8 $\pm$2.2 978
Table 5: All samples are divided into 3 complexity levels. We compute Pass@1.
Methods Single Multiple Complex Overall
Direct-Code 42.4% 29.0% 1.6% 28.5%
ReAct 56.3% 48.6% 14.6% 44.1%
Reflexion 67.8% 55.1% 48.1% 59.0%
Multi-agent 63.2% 64.7% 35.5% 57.4%
Data-Copilot 71.8% 70.0% 67.1% 70.2%
4.2 Comparison Results
Data-Copilot Significantly Reduces the Risk of Failure When Real-time Deployment. As demonstrated in Table 3, Compared to directly generating code, Data-Copilot achieves significant improvements across all LLMs, with particularly strong gains of +41.7, +18.7, and +29.2 on GPT-3.5-turbo, Llama3-70B, and Codellama-13B, respectively. By decoupling direct code generation into two stages: interface design in advance and real-time workflow deployment, Data-Copilot can elevate many smaller LLMs (e.g., Lllama3-70b, GPT-3.5) to GPT-4 level during real-time deployment. Through careful comparison, we observe that these pre-designed interfaces significantly reduce the occurrence of errors during real-time responses, as they have already been optimized and validated during the design phase. In contrast, directly generating code from scratch makes it easy to overlook certain details or make mistakes. Besides, we also found that even with a powerful LLM like GPT-4, Data-Copilot also achieves an improvement of over +4%, validating its generalizability.
Data-Copilot Outperforms Advanced Agent Strategies in Both Accuracy and Efficiency. As shown in Table 5, Data-Copilot surpasses all baseline strategies in success rate (Pass@1), and efficiency (#token). Compared to the best baseline (Reflexion), Data-Copilot achieved an +11.2% improvement in accuracy and a -75% reduction in token consumption. Besides, as shown in the last two rows of Table 5, Data-Copilot can seamlessly integrate with agent strategies such as ReAct and Reflexion into the workflow deployment. For example, when combined with ReAct, Data-Copilot invokes an interface, obtains an intermediate result, and then reasons to invoke the next interface, further improving its performance by +1.3.
Data-Copilot Reduces Repetitive Generation in the Real World. In the real world, numerous requests are similar or even repetitive. Data-Copilot can save 70% of the baseline’s consumption because its output is very concise and only contains interface names and arguments. In contrast, baseline strategies have to repetitively generate complete code for each request, consuming many more tokens.
Table 6: Different LLM combinations for two stages. Using a weaker LLM for interface design affects both the interface’s quantity and quality.
Direct Code Generation: GPT-3.5: 28.3, GPT-4: 70
LLM for Workflow Deployment
GPT-3.5 GPT-4
LLM for Interface Design GPT-3.5 31.9 49.6
GPT-4 70.2 74.6
Table 7: We ablated 3 designs from Data-Copilot.
Methods Pass@1(%)
Direct Code Generation 28.5
Data-Copilot 70.2
$w/o$ Self-Exploration Request 35.9
$w/o$ Interface Optimization 42.1
$w/o$ Interface-Code Hybrid 63.8
4.3 Analysis: Why Data-Copilot Can Bring Improvements?
Data-Copilot Exhibits a Pronounced Advantage on More Complex Scenarios. To better analyze the behavior of Data-Copilot, we categorize each test request into different subsets based on the number of entities it involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations). The statistics for each subset are shown in Section B.1. As shown in Table 5, for the single entity subset (denoted as Single), which represents the simplest case, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging multi-entities (denoted as Multiple) or multi-entities with complex relation (denoted as Complex), our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that in these complex cases, baseline methods often generate logically incorrect code on the first attempt and may omit some critical steps. In contrast, Data-Copilot only needs to invoke versatile interfaces and deploy workflows, significantly reducing the complexity of real-time responses. These findings highlight the advantages of Data-Copilot, which reduces task complexity by pre-designing interfaces in advance.
Interface Quality is the Key Factor. We analyze the effects of using different LLM combinations for two stages: interface design and workflow deployment. As shown in Table 7, we observed that when using a weaker LLM for interface design, the effectiveness is significantly diminished. For example, using GPT-3.5-turbo (1st stage) and GPT-4-turbo (2nd stage) resulted in a score of only 49.6, which is even lower than directly using GPT-4-turbo for code generation (70.2). This phenomenon can also be seen in other combinations. After manually checking the interfaces in this case, we found these interfaces designed by GPT-3.5-turbo are limited in number and of poor quality. Therefore, we chose to use GPT-4-turbo for interface design and optimization, while employing GPT-3.5-turbo for real-time deployment, achieving a balance between accuracy and efficiency.
4.4 Ablation Study
As shown in Table 7, we ablate Data-Copilot from the following three aspects:
  1. 1.
    Self-exploration Request. We ablate the self-exploration request process Section 2.1. Instead, we directly used seed requests for interface design and optimization. In this case, the performance dropped significantly by 34.3. This may be because there are too few seed requests, which are insufficient to design universe interfaces.
  2. 2.
    Interface Optimization. We ablate the interface optimization phase during interface design. It means every successfully designed interface was retained in the interface library. We observed that performance was also significantly affected (-28 points). Without interface optimization, there are many similar interfaces, which could hurt the effect of workflow invocation.
  3. 3.
    Interface-Code Hybrid Generation. Data-Copilot only can invoke interfaces during workflow deployment. Directly generating code is not allowed even when encountering requests that cannot be resolved by interfaces. After ablation, we observe a performance drop of 6.4 points. These losses are mainly caused by new requests that cannot be covered by existing interfaces. It highlights the flexibility of Data-Copilot, which can flexibly solve familiar requests by invoking interfaces or handle new requests by generating code directly.
5 Conclusion
We propose a data analysis agent, Data-Copilot, to address a wide array of data-related tasks. Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests. It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving. Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
Appendix
Appendix A Preprocessing of Workflow Deployment